# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

### v2

@@ Abstract
We propose a general-purpose approach for improving the ability of large language models (LLMs) to intelligently and adaptively gather information from a user or other external source using the framework of sequential Bayesian experimental design (BED). This enables LLMs to act as effective multi-turn conversational agents and interactively interface with external environments. Our approach, which we call BED-LLM (Bayesian experimental design with large language models), is based on iteratively choosing questions or queries that maximize the expected information gain (EIG) about the task of interest given the responses gathered previously. We show how this EIG can be formulated (and then estimated) in a principled way using a probabilistic model derived from the LLM’s predictive distributions and provide detailed insights into key decisions in its construction and updating procedure. We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20 Questions game and using the LLM to actively infer user preferences, compared to direct prompting of the LLM and other adaptive design strategies.
@@ 1 Introduction
Intelligent information gathering—the ability to ask the right questions at the right time—is fundamental to effective AI systems. However, despite their many successes, LLMs currently fall short on a crucial aspect of interactive intelligence: proactively seeking out information from a user or external environment in an intelligent and adaptive manner (Laban et al., 2025; Li et al., 2025c). For example, they have been shown to perform poorly on problems such as multi-turn guessing games (Bertolazzi et al., 2023; Zhang et al., 2024), task clarification (Chi et al., 2024), IT task automation (Jha et al., 2025), and multi-step tool use (Patil et al., 2025). In particular, while modern LLMs are often capable of producing coherent and insightful questions (or other external queries) in a single-turn setting, they typically struggle to appropriately tailor their questions to previously gathered responses on interactive tasks (Bertolazzi et al., 2023; Patil et al., 2025).
@@
There is, therefore, a pressing need to improve the ability of LLMs to adaptively ask questions based on previous responses, and gather information in a targeted manner. Such capabilities are essential for a wide variety of problems, such as clarifying user intent, personalizing model behavior to a particular user, or generally acting as effective multi-turn conversational agents. They are also critical if we want to use LLMs in data gathering tasks or as automated agents in decision-making pipelines (Wu et al., 2025). In turn, these capabilities are essential across domains ranging from medical diagnosis (Hirosawa et al., 2024), troubleshooting (Jha et al., 2025), preference learning (Handa et al., 2024; Chakraborty et al., 2024; Ouyang et al., 2022), and tutoring systems (Kestin et al., 2025; Liu et al., 2024a), to conducting automated surveys (Aher et al., 2023; Lee et al., 2024; Jacobsen et al., 2025), and AI–driven scientific inquiries (Lu et al., 2024; Mandal et al., 2025). Note that in all these problems it is not enough for the LLM to generate full sets of suitable questions up front, we need it to be able to adaptively choose questions that are tailored to the already-collected user responses.
@@
We propose to address this challenge using the framework of sequential Bayesian experimental design (BED, Lindley (1956); MacKay (1992); Chaloner & Verdinelli (1995); Sebastiani & Wynn (2000); Rainforth et al. (2024)), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment. Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem with a model derived from the LLM, wherein we iterate between choosing queries based on maximizing their expected information gain (EIG) and updating our beliefs with the information from the received response.
@@
We call our approach BED-LLM and show how its success is critically dependent on our precise model formulation, belief updating procedure, and EIG estimation strategy. In particular, we show that it is essential to formulate the model with a precise distribution pairing that does not solely rely on in-context learning to update beliefs and uses the LLM’s uncertainties in the space of answers rather than the more complicated underlying hypothesis space we are trying to learn in.
@@
Together, we find that these innovations provide substantial performance benefits over directly generating queries from the LLM and more basic approximations of the sequential BED framework. Specifically, we first find that BED-LLM provides substantial improvements in the success rate for the 20 Questions problem across a variety of LLMs and target quantities. For example, when guessing celebrities with a small model, Llama-3.1-8B, we observe a 5.8x gain in success rate. Second, we demonstrate noticeable improvements in using the LLM for movie recommendations, showing that these benefits hold even when the LLM’s predictive model differs from that of the answerer.
@@ 2 Problem Formulation and Background
There are two natural ways to improve LLMs’ ability to gather information: modifying the model itself (e.g. via test-time- or post-training) or altering how the model is used at deployment time. We focus on the latter, since information–gathering tasks rarely provide task-specific data upfront (e.g. a user’s unknown preferences), and deployment–time methods avoid the cost and difficulty of finetuning an LLM altogether and are applicable to any existing LLM. However, we emphasize that improvements at the model level (e.g., Zhang et al., 2024) would be complementary to our approach.
To formalize the notion of information gathering, we need a concrete idea of what we wish to learn about. We denote the target quantity of interest as $\theta$, which may represent, for example, a user’s preferences, the answer to a question, or a desired piece of content. We start with incomplete information about $\theta$, as represented by an initial belief distribution or prior, $p(\theta)$, but can refine these beliefs by making queries, $x\in\mathcal{X}$, to the user or some other external agent and receiving responses, $y\in\mathcal{Y}$, that are informative about $\theta$. Multiple such queries, $x_{1},\dots,x_{n}$, can be adaptively selected in a sequential decision-making process where we iteratively choose each $x_{t}$ based on the collected history $h_{t-1}:=(x_{i},y_{i})_{i=1:t-1}$. As our history grows, we will update our belief distribution to obtain $p(\theta;h_{t-1})$ via some model updating procedure. [^1] We carefully distinguish between explicit probabilistic conditioning, i.e. $p(a|b)$, and more general dependency, $p(a;b)$. The former corresponds to the conditional distribution of an associated joint distribution, $p(a,b)$, while the latter may not. Here, $h_{t-1}$ influences our distribution on $\theta$, but it is not derived via a joint distribution. In the LLM setting, there is considerable flexibility in how $p(\theta;h_{t-1})$ is constructed, as discussed in §3.3 and §4. While $p(\theta;h_{t-1})$ need not be explicitly defined, it provides the foundation for our information-theoretic method of query selection.
@@
2.1 In–Context Updating of the belief distribution
A natural and cheap way to incorporate the interaction history into the LLM is to include it in the context (Brown et al., 2020). If the LLM’s distribution over generated text, $z\!\in\!\mathcal{Z}$, is $p_{\mathrm{LLM}}(z)$ given appropriate prompting, then $p_{\mathrm{LLM}}(z;h_{t-1})$ is an updated distribution with the previous question–response pairs in context. From this, we can derive an updated belief distribution over $\theta$. Most simply, this can be done by using $p_{\mathrm{LLM}}(z;h_{t-1})$ to directly query about $\theta$ (e.g. if $\theta$ is some preference, we could prompt the LLM to predict this preference). However, as we show later, this approach often fails to appropriately incorporate the information from $h_{t-1}$, leading to a belief distribution inconsistent with past observations. This is consistent with recent work that shows that in context updating does not treat all contextual information equally (Kossen et al., 2024; Liu et al., 2024b; Zhang et al., 2024). In §3.3, we introduce a more robust method for deriving $p(\theta;h_{t-1})$.
2.2 Information–Theoretic Experimental Design
The core of the BED framework is a joint generative model $p(\theta,y;x)$ over the target quantity $\theta$ and outcomes, $y$, given designs $x$. Most commonly, this is specified as a Bayesian model using a prior $p(\theta)$ and likelihood $p(y|\theta;x)$. In the general case, designs are then chosen to maximize the expectation of some utility function $U(\theta,y,x)$ under this model: we choose $x^{*}=\mathrm{argmax}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$. The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p(\theta)\log p(y;x)$, where $p(\theta)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design $x$. This leads to an objective corresponding to the expected information gain (EIG) in $\theta$ (Lindley, 1956; 1972),
$\text{EIG}_{\theta}(x)$ $=\mathrm{H}[p(\theta)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (1)
@@ 2.2 Information–Theoretic Experimental Design
where $\mathrm{H}$ denotes the Shannon entropy (i.e., $\mathrm{H}[p(\theta)]=-\mathbb{E}_{p(\theta)}[\log p(\theta)]$). We can thus equivalently think of the EIG as: a) the mutual information between $\theta$ and $y$, b) the expected information gain over possible data simulated from our model (where the information gain is defined as the reduction in entropy from our prior on $\theta$ to the posterior), or c) the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn, 2000).
Working with the EIG is highly suited to a sequential or adaptive design approach, wherein it is generally referred to as sequential BED or Bayesian Adaptive Design (Rainforth et al., 2024). Because the EIG is only a function of our underlying model, when we update the model as new data becomes available, our EIG design objective will naturally update as well. Specifically, to derive the incremental EIG (Cavagnaro et al., 2010) for the $t$-th query, $\text{EIG}_{\theta}(x_{t};h_{t-1})$, we simply replace the joint $p(\theta,y;x)$ in the above formulation with the updated joint $p(\theta,y_{t};h_{t-1},x_{t})$, with all marginals a conditionals derived from this (e.g. $p(y;x)$ becomes $p(y_{t};h_{t-1},x_{t})$). Here this updated joint conventionally comes from a Bayesian update of the original model. However, in many cases, this is not practical and other non-Bayesian updates are performed instead, e.g. in active learning the update often actually corresponds to retraining the model with the new data (Gal et al., 2017; Bickford Smith et al., 2023).
3 Sequential Bayesian Experimental Design for LLMs
The sequential BED framework described in §2.2 requires two core components to be specified by the user: a) an initial joint model $p(\theta,y;x)$ over hypotheses $\theta$ and outcomes $y$, given chosen experiment $x$, and b) a procedure to derive an updated model $p(\theta,y_{t};h_{t-1},x_{t})$ after observing $h_{t-1}$. In the LLM setting, there is significant flexibility in these critical design decisions. In particular, there are many ways to derive a suitable joint distribution from the LLM and its ability to learn in-context provides opportunities for update methods that go beyond standard Bayesian model updates.
@@ 3 Sequential Bayesian Experimental Design for LLMs
A major challenge in the LLM setting is that unlike conventional probabilistic models, in general, $p_{\mathrm{LLM}}(\theta)\,p_{\mathrm{LLM}}(y;[\theta,x])\neq p_{\mathrm{LLM}}(y;x)\,p_{\mathrm{LLM}}(\theta;[x,y])$. That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing), than if we first sample $y$ then sample $\theta$ with $y$ in context (data–estimation pairing). Moreover, we can deviate from the distribution directly induced by the LLM on one or both variables. The success of using BED with LLMs turns out to be critically dependent on these choices.
We delay proper discussion of this complex issue until §4, where we will see that the preferable setup can depend on problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$. For now, we will focus on using the prior–likelihood pairing; we will argue in §4 that this is the advantageous setup in many practical scenarios. While we will generally use the LLM’s directly induced distribution for the likelihood, we allow the prior to deviate from this in a problem–specific manner. As such, our initial joint model will be $p(\theta,y;x)=p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])$.
@@
Optimally updating the joint model in this setting requires incorporating new observations in a way that both fully captures the information from new data and is computationally tractable. At one extreme, we could perform full Bayesian updates via approximate inference, as in classical sequential BED. However, this demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static likelihood queries. At the other extreme, simple in-context updating, $p(\theta;h_{t-1})=p_{\mathrm{LLM}}(\theta;h_{t-1})$, is cheap but, as we show later, fails to reliably capture information from new data, leading to inconsistent belief states and undermining the sequential BED approach. As we discuss in §3.3, we therefore employ a strategy that is somewhere between the two: drawing samples in a way that utilizes $p_{\mathrm{LLM}}(\theta;h_{t-1})$ while encouraging diversity, then filtering out samples that are actually not compatible with $h_{t-1}$ and renormalizing. We refer to the resulting distribution as $p_{f}(\theta;h_{t-1})$. We do not update our likelihood model $p_{\mathrm{LLM}}(\theta;h_{t-1})$; see Section˜A.1 for a discussion.
@@
We now introduce our specific algorithmic approach, BED-LLM. Here, the queries will correspond to our designs, $x$ (assumed to be in form of questions posed to the user in the following for simplicity, but could also be, e.g., external function calls, document retrieval, web search, etc), and the responses received will correspond to our outcomes, $y$. Using the LLM to derive joint models over these outcomes and the target variables $\theta$ given histories, $h_{t-1}$ as described above, we can interleave choosing informative questions by optimizing the incremental EIG, $\text{EIG}_{\theta}(x_{t};h_{t-1})$, and updating our underlying model based on the received question–response pairs. Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn:
  1. 1.
@@
    Generate candidate questions3.1): Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\text{cand}}$, by appropriate sampling of the LLM based on the conversational context $h_{t-1}$.
  2. 2.
@@
    Compute EIG estimator3.2): For each candidate $x_{t}\in\mathcal{X}^{\text{cand}}$, estimate $\text{EIG}_{\theta}(x_{t};h_{t-1})$.
  3. 3.
    Select and ask optimal question: Choose the question $x_{t}\in\mathcal{X}^{\text{cand}}$ that yields the highest estimated EIG. Pose $x_{t}$ to the user, observe response $y_{t}$, and update the history, $h_{t}=(h_{t-1},(x_{t},y_{t}))$.
  4. 4.
    Construct updated joint (Section˜3.3): $p(\theta,y_{t+1};h_{t},x_{t+1})=p_{f}(\theta;h_{t})p_{\mathrm{LLM}}(y_{t+1};[\theta,x_{t+1}])$, using the new history and return to Step 1 (unless a termination criterion has been achieved).
Note our belief state on $\theta$ after the $t$-th turn is simply given by $p_{f}(\theta;h_{t})$.

### v3

@@ Abstract
We propose a general-purpose approach for improving the ability of large language models (LLMs) to intelligently and adaptively gather information from a user or other external source using the framework of sequential Bayesian experimental design (BED). This enables LLMs to act as effective multi-turn conversational agents and interactively interface with external environments. Our approach, which we call BED-LLM (Bayesian experimental design with large language models), is based on iteratively choosing questions or queries that maximize the expected information gain (EIG) with respect to a variable of interest given the responses gathered previously. We show how this EIG can be formulated (and then estimated) in a principled way using a probabilistic model derived from the LLM’s predictive distributions and provide detailed insights into key decisions in its construction and updating procedure. We find that BED-LLM achieves substantial gains in performance across a wide range of tests based on the 20 Questions game and using the LLM to actively infer user preferences, compared to purely prompting-based design generation and other adaptive design strategies.
@@ 1 Introduction
Intelligent information gathering—the ability to ask the right questions at the right time—is fundamental to effective AI systems. However, despite their many successes, LLMs currently fall short on proactively seeking out information from a user or external environment in an adaptive manner (Laban et al., 2025; Li et al., 2025c). For example, they have been shown to perform poorly on multi-turn guessing games (Bertolazzi et al., 2023; Zhang et al., 2024), task clarification (Chi et al., 2024), IT task automation (Jha et al., 2025) and multi-step tool use (Patil et al., 2025). In particular, while modern LLMs are often capable of producing coherent and insightful questions (or other external queries) in a single-turn setting, they typically struggle to appropriately tailor their questions to previously gathered responses on interactive tasks (Bertolazzi et al., 2023; Patil et al., 2025).
@@
There is, therefore, a pressing need to improve the ability of LLMs to adaptively ask questions based on previous responses, and gather information in a targeted manner. Such capabilities are essential for a wide variety of problems, such as clarifying user intent, personalizing model behavior to a particular user, or generally acting as effective multi-turn conversational agents. They are also critical if we want to use LLMs in data gathering tasks or as automated agents in decision-making pipelines (Wu et al., 2025). In turn, these capabilities are essential across domains ranging from medical diagnosis (Hirosawa et al., 2024), troubleshooting (Jha et al., 2025), preference learning (Chakraborty et al., 2024; Handa et al., 2024; Ouyang et al., 2022) and tutoring systems (Kestin et al., 2025; Liu et al., 2024a), to conducting automated surveys (Aher et al., 2023; Jacobsen et al., 2025; Lee et al., 2024) and AI-driven scientific inquiries (Lu et al., 2024; Mandal et al., 2025). Note that in all these problems it is not enough for the LLM to generate full sets of suitable questions up front: we need it to be able to adaptively choose questions that are tailored to the already-collected user responses.
@@
We propose to address this challenge using the framework of sequential Bayesian experimental design (BED; Chaloner & Verdinelli, 1995; Lindley, 1956; MacKay, 1992; Rainforth et al., 2024; Sebastiani & Wynn, 2000), which provides a model-based, information-theoretic mechanism for making adaptive design decisions, given a generative model of the experiment. Specifically, we show how the problem of interactive information gathering with LLMs can be formulated as a sequential experimental design problem with a model derived from the LLM, wherein we iterate between choosing queries by maximizing expected information gain (EIG) in a variable of interest and updating our beliefs with the information from the received response.
@@
We call our approach BED-LLM and show how its success is critically dependent on our precise model formulation, belief updating procedure and EIG estimation strategy. In particular, we show that it is essential to formulate the model with a precise distribution pairing that does not solely rely on in-context learning to update beliefs and uses the LLM’s uncertainties in the space of answers rather than the more complicated underlying hypothesis space we are trying to learn in.
@@
Together, we find that these innovations provide substantial performance benefits over directly generating queries from the LLM and more basic approximations of the sequential BED framework. Specifically, we first find that BED-LLM provides substantial improvements in the success rate for the 20 Questions problem across a variety of LLMs and target quantities. On average, BED-LLM improves the success rate by 37.4 percentage points compared to direct prompting of the LLM, with the success rate more than doubling in over half of the setups considered and never decreasing. Second, we demonstrate noticeable improvements in using the LLM for movie recommendations, showing that these benefits hold even when the LLM’s predictive model differs from that of the answerer.
@@ 2 Problem formulation and background
There are two natural ways to improve LLMs’ ability to gather information: modifying the model itself (e.g. via finetuning) or altering how the model is used at deployment time. We focus on the latter, since information-gathering tasks rarely provide task-specific data upfront (e.g. a user’s unknown preferences), and deployment-time methods avoid the cost and difficulty of finetuning an LLM altogether and are applicable to any existing LLM. However, we emphasize that improvements at the model level (e.g. Zhang et al., 2024) would be complementary to our approach.
To formalize the notion of information gathering, we need a concrete idea of what we wish to learn about. We denote the target quantity of interest as $\theta$, which may represent, for example, a user’s preferences, the answer to a question, or a desired piece of content. We start with incomplete information about $\theta$, as represented by an initial belief distribution or prior, $p(\theta)$, but can refine these beliefs by making queries, $x\in\mathcal{X}$, to the user or some other external agent and receiving responses, $y\in\mathcal{Y}$, that are informative about $\theta$. Multiple such queries, $x_{1},\dots,x_{T}$, can be adaptively selected in a sequential decision-making process where we iteratively choose each $x_{t}$ based on the collected history, $h_{t-1}:=(x_{i},y_{i})_{i=1}^{t-1}$. As our history grows, we will update our belief distribution to obtain $p(\theta;h_{t-1})$ via some model updating procedure. [^1] We carefully distinguish between explicit probabilistic conditioning, i.e. $p(a|b)$, and more general dependency, $p(a;b)$. The former corresponds to the conditional distribution of an associated joint distribution, $p(a,b)$, while the latter may not. Here, $h_{t-1}$ influences our distribution on $\theta$, but it is not derived via a joint distribution. In the LLM setting, there is considerable flexibility in how $p(\theta;h_{t-1})$ is constructed, as discussed in Section˜3.1 and Section˜4. While $p(\theta;h_{t-1})$ need not be explicitly defined, it provides the foundation for our information-theoretic method of query selection.
@@
2.1 In-context updating of beliefs
A natural and cheap way to incorporate the interaction history into the LLM is to include it in the context (Brown et al., 2020). If the LLM’s distribution over generated text, $z\in\mathcal{Z}$, is $p_{\mathrm{LLM}}(z)$ given appropriate prompting, then $p_{\mathrm{LLM}}(z;h_{t-1})$ is an updated distribution with the previous question–response pairs in context. From this, we can derive an updated belief distribution over $\theta$. Most simply, this can be done by using $p_{\mathrm{LLM}}(z;h_{t-1})$ to directly query about $\theta$ (e.g. if $\theta$ is some preference, we could prompt the LLM to predict this preference). However, as we show later, this approach often fails to appropriately incorporate the information from $h_{t-1}$, leading to a belief distribution inconsistent with past observations. This is consistent with recent work that shows that in context updating does not treat all contextual information equally (Kossen et al., 2024; Liu et al., 2024b; Zhang et al., 2024). In Section˜3.1, we introduce a more robust method for deriving $p(\theta;h_{t-1})$.
2.2 Information-theoretic experimental design
The core of the BED framework is a joint generative model, $p(\theta,y;x)$, over the target quantity, $\theta$, and the outcome, $y$, corresponding to a design, $x$. Typically this is specified as a Bayesian model using a prior, $p(\theta)$, and a likelihood, $p(y|\theta;x)$. In the general case, designs are chosen to maximize the expectation of some utility function, $U(\theta,y,x)$, under this model: we choose $x^{*}=\mathop{\mathrm{arg\,max}}_{x}\mathbb{E}_{p(\theta,y;x)}\left[U(\theta,y,x)\right]$. The most common choice is to take $U(\theta,y,x)=\log p(\theta,y;x)-\log p(\theta)\log p(y;x)$, where $p(\theta)$ and $p(y;x)$ are the marginal distributions on $\theta$ and $y$ implied by our joint model and we have assumed that our current beliefs on $\theta$ are independent of the design, $x$. This leads to an objective corresponding to the expected information gain (EIG) in $\theta$ (Lindley, 1956; 1972),
$\mathrm{EIG}_{\theta}(x)$ $=\mathrm{H}[p(\theta)]-\mathbb{E}_{p(y;x)}[\mathrm{H}[p(\theta|y;x)]]$ (1)
@@ 2.2 Information-theoretic experimental design
where $\mathrm{H}$ denotes the Shannon entropy (i.e. $\mathrm{H}[p(\theta)]=-\mathbb{E}_{p(\theta)}[\log p(\theta)]$). We can thus equivalently think of the EIG as (a) the mutual information between $\theta$ and $y$, (b) the expected reduction in entropy over $\theta$ (i.e. information gain in $\theta$) from observing data simulated from our model, or (c) the expected reduction in entropy over data from observing $\theta$ simulated from our prior (Sebastiani & Wynn, 2000).
Working with the EIG is highly suited to a sequential or adaptive design approach, generally referred to as sequential BED or Bayesian adaptive design (Rainforth et al., 2024). Because the EIG is only a function of our underlying model, when we update the model as new data becomes available, our EIG design objective will naturally update as well. Specifically, to derive the incremental EIG (Cavagnaro et al., 2010) for the $t$-th query, $\mathrm{EIG}_{\theta}(x_{t};h_{t-1})$, we simply replace the joint in the above formulation, $p(\theta,y;x)$, with the updated joint, $p(\theta,y_{t};h_{t-1},x_{t})$, with all marginals and conditionals derived from this (e.g. $p(y;x)$ becomes $p(y_{t};h_{t-1},x_{t})$). Here this updated joint conventionally comes from a Bayesian update of the original model. However, in many cases, this is not practical and other non-Bayesian updates are performed instead. For example, in active learning the update often actually corresponds to retraining the model with the new data (Bickford Smith et al., 2023; Gal et al., 2017).
3 Sequential Bayesian experimental design with LLMs
The sequential BED framework described in Section˜2.2 requires two core components to be specified by the user: (a) an initial joint model, $p(\theta,y;x)$, over hypotheses, $\theta$, and outcomes, $y$, given a design, $x$; and (b) a procedure to derive an updated model, $p(\theta,y_{t};h_{t-1},x_{t})$, after observing $h_{t-1}$. In the LLM setting, there is significant flexibility in these critical methodological decisions. In particular, there are many ways to derive a suitable joint distribution from the LLM and its ability to learn in-context provides opportunities for update methods that go beyond standard Bayesian model updates.
@@ 3 Sequential Bayesian experimental design with LLMs
A major challenge in the LLM setting is that unlike conventional probabilistic models, in general, $p_{\mathrm{LLM}}(\theta)\,p_{\mathrm{LLM}}(y;[\theta,x])\neq p_{\mathrm{LLM}}(y;x)\,p_{\mathrm{LLM}}(\theta;[x,y])$. That is, we induce a different joint distribution if we first sample $\theta$ then sample $y$ with $\theta$ in context (which we refer to as the prior–likelihood pairing) versus if we first sample $y$ then sample $\theta$ with $y$ in context (data–estimation pairing). Moreover, we can deviate from the distribution directly induced by the LLM on one or both variables. The success of using BED with LLMs turns out to be critically dependent on these choices.
We delay proper discussion of this complex issue until Section˜4, where we will see that the preferable setup can depend on problem setting and, in particular, the relative complexity of spaces of $\theta$ and $y$. For now, we will focus on using the prior–likelihood pairing; we will argue in Section˜4 that this is the advantageous setup in many practical scenarios. While we will generally use the LLM’s directly induced distribution for the likelihood, we allow the prior to deviate from this in a problem-specific manner. As such, our initial joint model will be $p(\theta,y;x)=p(\theta)p_{\mathrm{LLM}}(y;[\theta,x])$.
@@
Optimally updating the joint model in this setting requires incorporating new observations in a way that both fully captures the information from new data and is computationally tractable. At one extreme, we could target full Bayesian updates via approximate inference, as in classical sequential BED. However, this demands a prohibitively large number of LLM evaluations to accurately approximate the posterior, and it does not exploit the power of the LLM as a probabilistic generative model, where autoregressive sequential rollouts often lead to more nuanced and diverse behavior than repeated static queries. At the other extreme, simple in-context updating, with $p(\theta;h_{t-1})=p_{\mathrm{LLM}}(\theta;h_{t-1})$, is cheap but, as we show later, fails to reliably capture information from new data, leading to inconsistent belief states and undermining the sequential BED approach. As we discuss in Section˜3.1, we therefore employ a strategy that is somewhere between the two: drawing samples in a way that utilizes $p_{\mathrm{LLM}}(\theta;h_{t-1})$ while encouraging diversity, then filtering out samples that are actually not compatible with $h_{t-1}$ and renormalizing. We refer to the resulting distribution as $p_{f}(\theta;h_{t-1})$. We do not update our likelihood model, $p_{\mathrm{LLM}}(\theta;h_{t-1})$; see Section˜F.3 for empirical comparison of updated vs. static likelihoods in our experiments, and a discussion.
@@
We now introduce our specific algorithmic approach, BED-LLM. Here, the queries will correspond to our designs, $x$ (assumed to be in form of questions posed to the user in the following for simplicity, but could also be external function calling, document retrieval, web search, etc), and the responses received will correspond to our outcomes, $y$. Using the LLM to derive joint models over these outcomes and the target variable, $\theta$, given the history, $h_{t-1}$, as described above, we can interleave choosing informative questions by optimizing the incremental EIG, $\mathrm{EIG}_{\theta}(x_{t};h_{t-1})$, and updating our underlying model based on the received question–response pairs. Specifically, BED-LLM iterates over the following key steps, where $t$ indexes the current turn (see also Figure˜1).
  1. (A)
    Extract beliefs (Section˜3.1). Use joint model $p(\theta,y_{t};h_{t-1},x_{t})=p(\theta)p_{\mathrm{LLM}}(y_{t};[\theta,x_{t}])$, where $h_{0}=\emptyset$, for $t=1$ and $p(\theta,y_{t};h_{t-1},x_{t})=p_{f}(\theta;h_{t-1})p_{\mathrm{LLM}}(y_{t};[\theta,x_{t}])$ for $t>1$.
  2. (B)
@@
    Generate candidate questions (Section˜3.2). Propose a candidate set of $M$ diverse, multiple-choice questions, $\mathcal{X}^{\mathrm{cand}}$, by appropriate sampling of the LLM based on the history, $h_{t-1}$.
  3. (C)
@@
    Estimate EIG (Section˜3.3). For each candidate, $x_{t}\in\mathcal{X}^{\mathrm{cand}}$, estimate $\mathrm{EIG}_{\theta}(x_{t};h_{t-1})$.
  4. (D)
    Select and ask the best question. Choose $x_{t}$ by EIG maximization and pose to the user.
  5. (E)
    Update the history. Observe response, $y_{t}$, and update the history to $h_{t}=(h_{t-1},(x_{t},y_{t}))$.
Current history: $h_{2}=((\text{``Born in the 20th century?''},\text{``Yes''}),(\text{``Is this person male?''},\text{``Yes''}))$. (A) Extract beliefs (Section˜3.1). Construct hypothesis set $\Theta^{\mathrm{cand}}=\{\theta_{n}\}_{n=1}^{N}$ by sampling candidate hypotheses, $\theta\sim p_{\mathrm{LLM}}(\theta;h_{2})$, and rejecting any $\theta$ that is inconsistent with $h_{2}$. Use joint model $p(\theta,y;h_{2},x)=p_{f}(\theta;h_{2})p_{\mathrm{LLM}}(y;[\theta,x])$, where $p_{f}(\theta;h_{2})$ is uniform over $\Theta^{\mathrm{cand}}$. Barack Obama Steve Irwin Hugh Laurie Banksy Elvis Presley (B) Generate candidate questions (Section˜3.2). Sample $\tilde{x}_{1:M}\sim p_{\mathrm{LLM}}(\tilde{x}_{1:M};[h_{t-1},\Theta^{\mathrm{cand}}])$. $\tilde{x}_{1}$: “Was this person born in Antarctica?” $\tilde{x}_{2}$: “Was this person born in the 19th century?” $\tilde{x}_{3}$: “Is this person an artist?” $\tilde{x}_{4}$: “Is this person European?” $\tilde{x}_{5}$: “Does this person prefer thrash metal over death metal?” (C) Estimate EIG (Section˜3.3). Compute $\underbrace{\mathrm{H}[\hat{p}(y;[h_{2},x])]}_{\text{marginal entropy}}-\underbrace{\tfrac{1}{N}\textstyle\sum_{n=1}^{N}\mathrm{H}[p_{\mathrm{LLM}}(y;[\theta_{n},x])]}_{\text{expected conditional entropy}}\approx\mathrm{EIG}_{\theta}(x;h_{2})$. Question Marg. Exp. cond. EIG Intuition $\tilde{x}_{1}$ Born in Antarctica? $\approx 0$ $\approx 0$ $\approx 0$ Answer “No” for all $\theta_{n}$ $\tilde{x}_{2}$ Born in 19th century? $\approx 0$ $\approx 0$ $\approx 0$ Redundant given $h_{2}$ $\tilde{x}_{3}$ An artist? $0.56$ $0.41$ $0.15$ Uneven split $\tilde{x}_{4}$ European? $0.97$ $0.17$ $0.80$ Balanced split; crisp answers $\tilde{x}_{5}$ Thrash vs. death metal? $0.89$ $0.88$ $0.01$ Uncertain even given $\theta$ (D) Select and ask the best question. Choose $x_{3}=\tilde{x}_{4}$ by maximizing EIG. Output this to the user. (E) Update the history. Observe answer, $y_{3}$. Set $h_{3}=(h_{2},(x_{3},y_{3}))$.
Figure 1: BED-LLM applied to the 20 Questions game involves repeatedly constructing a belief state, generating candidate questions, estimating and maximizing EIG to select a question, and interacting with the user to gather new data. The contrast between $\tilde{x}_{4}$ and $\tilde{x}_{5}$ illustrates the benefit of using EIG with a non-deterministic likelihood: $\tilde{x}_{5}$ has high marginal entropy (answers are uncertain), but its expected conditional entropy is equally high (answers are uncertain even given $\theta$), so nearly nothing is expected to be learned, and $\tilde{x}_{4}$ should be favoured as a result. Numerical values are illustrative.
3.1 Prior construction and belief updating
The Savage axioms (Savage, 1954) tell us that a rational agent should update its beliefs in a Bayesian manner. However, doing full Bayesian updates to our model as the history grows is generally impractical for computational reasons in the LLM setting, as it requires approximate inference and this, in turn, typically requires large numbers of expensive likelihood evaluations. Furthermore, the Savage axioms only hold if our (implied) prior truly represents our beliefs, but we find that $p_{\mathrm{LLM}}(\theta)$ is typically heavily overconfident on a small number of possible hypotheses and can struggle to convey the full range of possibilities even with careful prompting and a high temperature (see Figure˜5).
To avoid these shortfalls, we instead propose an approach that balances tractability and faithfulness. Although we will still use $p_{\mathrm{LLM}}(\theta;h_{t-1})$ as the basis for deriving our belief state over $\theta$ (i.e. our intermediate prior), we make various alterations to effectively incorporate historical information and ensure diversity. Our derived distribution, which we refer to as $p_{f}(\theta;h_{t-1})$, differs from $p_{\mathrm{LLM}}(\theta;h_{t-1})$ in two key ways. First, we filter the generated hypotheses according to whether they are compatible with the history $h_{t-1}$. We do this by using the LLM to check the compatibility of each sampled $\theta$ with all the previous question–answer pairs in $h_{t-1}$ (using $p_{\mathrm{LLM}}(y_{i};[\theta,x_{i}])$ for $i\in(1,2,\ldots,t-1)$) and then rejecting that sample if an incompatibility is found. Specifically, a sample is rejected if the likelihood of an observed answer falls below a predefined threshold, chosen to balance robustness to model uncertainty against the need to enforce strict historical coherence. To reduce the computational cost of generating and evaluating hypotheses, we further include a hypothesis-retention mechanism: any hypotheses from the previous turn which remain consistent with the most recent question and observation are retained in the hypothesis set without regeneration.
Second, we make a number of modifications to promote diversity. Rather than generate candidates independently, we prompt the LLM to generate batches of candidates using a prompt encouraging diversity. After filtering these candidates as above and removing duplicates, we then impose a uniform distribution. Details of our exact setup for $p_{f}(\theta;h_{t-1})$ are given in Appendix˜E.